OpenStudy (vortish):

find the term that should be added to the expression to form a perfect square trinomial. Write the perfect square trinomial in factored form x^2+5x

jimthompson5910 (jim_thompson5910):

In general to make x^2 + bx a perfect square, you add on (b/2)^2 = (b^2)/4

jimthompson5910 (jim_thompson5910):

That will give you x^2 + bx + (b^2)/4 which will factor to (x+b/2)^2

OpenStudy (vortish):

Mr jim that is so confusing to me can you break this down some so i understand it

jimthompson5910 (jim_thompson5910):

b = ??? in this case

OpenStudy (vortish):

yes please

jimthompson5910 (jim_thompson5910):

what is the value of b in this case

OpenStudy (vortish):

would that be five

jimthompson5910 (jim_thompson5910):

so (b/2)^2 = (5/2)^2 = 25/4

jimthompson5910 (jim_thompson5910):

which means this is the number you add on to make x^2 + 5x a perfect square

OpenStudy (vortish):

so the entire formula would be X^2+5x+25/4

jimthompson5910 (jim_thompson5910):

yes that's a perfect square

jimthompson5910 (jim_thompson5910):

which factors to....?

OpenStudy (vortish):

i know that sense five is a general number that is in the whole equation

jimthompson5910 (jim_thompson5910):

x^2 + bx + (b^2)/4 factors to (x+b/2)^2 b = 5, so x^2+5x+25/4 factors to (x+5/2)^2

OpenStudy (vortish):

ok that makes sense Mr Thompson thank you so much for your time and expertise

jimthompson5910 (jim_thompson5910):

sure thing